### 16 Mar 2026 v1

# Rigidity and Structural Asymmetry of Bounded Solutions

Abstract
In this manuscript, we introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter. We identify a Rotation number hypothesis on the non-homogeneous term, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\overline{s}$ lying in the critical strip, are both bounded on $1,+\infty)$, then $\Re(s)=\tfrac{1}{2}$.
Keywords: Non-homogeneous complex linear differential equation, bounded solutions, Volterra integral equation.
MSC subject classifications: 34A30, 34E05
1 Main result
In this manuscript, we study the following non-homogeneous linear complex differential equations:
$\dot{\theta}=wt^{-1}\theta+t^{-1}\eta(t),\quad\theta(1)=\frac{1}{1-w},\quad\theta:[1,+\infty)\to\mathbb{C},$ (1)
where $w\in\mathbb{C}$ such that $\Im(w)\neq 0$ is the parameter and the non-homogeneous term $\eta\in L^{\infty}(\mathbb{C})$ where $L^{\infty}(\mathbb{C})$ is the set of bounded and locally integrable functions from $1,+\infty)$ to $\mathbb{C}$. As formulated in the Dynamical Conjecture in [1], we assume that the function $\eta$ satisfies the following Rotation number hypothesis:
$\exists\rho\in\mathbb{C}^{*}:\quad\sup_{t\geq 1}\Big|\int_{1}^{t}\,\big(\eta(u)-\rho\big)du\Big|<+\infty,$ (H)
The complex number $\rho$ is called the rotation number of $\eta$.
We are interested in the initial conditions of the bounded solutions of the previous differential equation. To this end, we analyze the behavior of the transformed solution $t\mapsto(1-w)t^{-i\Im(w)}\theta(t)$, which leads us to consider the following differential equation:
$\dot{x}=\Re(w)t^{-1}x+(1-w)t^{-1-i\Im(w)}\eta(t),$ (2)
$t\in[1,+\infty),\quad x(1)=1,\quad x:[1,+\infty)\to\mathbb{C}.$
Since $\eta\in L^{\infty}(\mathbb{C})$, the function
$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1.$
is absolutely continuous on $1,+\infty)$. The differential equation (2) is a non-homogeneous linear differential equation. Then, there exists a unique continuous solution $\psi_{\eta,w}(\cdot)$ of (2) such that $\psi_{w}(1)=1$, which is given by
$\psi_{\eta,w}(t)=t^{\Re(w)}\Big[1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (3)
Let us introduce some notations.
Notation 1.
For every $\eta\in L^{\infty}(\mathbb{C})$ and $w\in\mathbb{C}$, we denote by $\psi_{\eta,w}(\cdot)$ the unique continuous solution of the differential equation (2) given by equation (3).
Notation 2.
Denote by $\mathbb{C}_{+}$ the right half-plane defined as
$\mathbb{C}_{+}:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
For every $\eta\in L^{\infty}(\mathbb{C})$, let $\mu_{\eta}$ denote the function $\mu_{\eta}:\mathbb{C}_{+}\to\mathbb{C}$, defined as
$\mu_{\eta}(w)=-1-(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall w\in\mathbb{C}_{+}.$
The function $\mu_{\eta}$ is defined for all $s\in\mathbb{C}_{+}$. Indeed, the integral is absolutely convergent, since $\eta\in L^{\infty}(\mathbb{C})$ and since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$. In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
Lemma 3.
Let $\eta\in L^{\infty}(\mathbb{C})$ and $w\in\mathbb{C}_{+}$. Consider the continuous solution $\psi_{\eta,w}(\cdot)$ of the differential equation (2), as introduced in Notation 1. Then,
$\sup_{t\geq 1}\left|\psi_{\eta,w}(t)\right|<+\infty\quad\text{if and only if}\quad\mu_{\eta}(w)=0,$
where the function $\mu_{\eta}$ is defined in Notation 2. Furthermore, if $\eta$ satisfies the rotation number hypothesis (H), then
$\sup_{t\geq 1}\left|t\left(\psi_{\eta,w}\big(t\big)-\rho\frac{w-1}{w}t^{-i\Im(w)}\right)\right|<+\infty,$
where $\rho$ denotes the rotation number of $\eta$.
Proof.
Let $w\in\mathbb{C}_{+}$ be fixed. Suppose that $\sup_{t\geq 1}|\psi_{\eta,w}(t)|<+\infty$. By equation (3), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
Since $\Re(w)>0$, we must have
$\Big|1+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|=0,$
which gives $\mu_{\eta}(w)=0$. Now, prove that $\sup_{t\geq 1}|\psi_{\eta,w}(t)|<+\infty$. By definition of $\mu_{\eta}(w)$, we have
$\mu_{\eta}(w)+1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du=-(1-w)\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\ \forall t\geq 1.$
Suppose that $\mu_{\eta}(w)=0$, using equation (3), we obtain
$\psi_{\eta,w}(t)=-(1-w)t^{\Re(w)}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (4)
By Hypothesis $\eta\in L^{\infty}(\mathbb{C})$ then there exists a constant $c>0$ such that $|\eta(t)|<c$ almost everywhere, we obtain
$\sup_{t\geq 1}|\psi_{\eta,w}(t)|\leq c\frac{|1-w|}{\Re(w)}<+\infty.$
Now, prove the second item of the present Lemma. Suppose that the function $\eta$ satisfies the hypothesis (H). Then there exists $\tilde{c}>0$ such that
$\Big|\int_{1}^{t}\,\big(\eta(u)-\rho\big)du\Big|<\tilde{c},\quad\forall t\geq 1,$
where $\rho$ is the rotation number of $\eta$. Hence, the equation (4) can be written as
$\psi_{\eta,w}(t)=\rho\frac{w-1}{w}t^{-i\,\Im(w)}-(1-w)t^{{\Re(w)}}\int_{t}^{+\infty}u^{-1-w}(\eta(u)-\rho)du,$
Using the integration by parts formula, we obtain
$\psi_{\eta,w}(t)=\rho\frac{w-1}{w}t^{-i\,\Im(w)}-(1-w^{2})t^{\Re(w)}\int_{t}^{+\infty}u^{-2-w}\int_{t}^{u}(\eta(v)-\rho)dvdu.$
For every $t\geq 1$, we get
$\left|\psi_{\eta,w}(t)-\rho\frac{w-1}{w}t^{-i\,\Im(w)}\right|<2\tilde{c}\frac{|1-w^{2}|}{1+\Re(w)}t^{-1}.$
We denote by $B\subset\mathbb{C}$ the critical strip, excluding the critical line and the real axis, defined as
$B:=\Big\{w\in\mathbb{C}:\Re(w)\in(0,1),\ \Re(w)\neq\tfrac{1}{2},\ \Im(w)\neq 0\Big\}.$
The following main result provides an answer to the Dynamical Conjecture formulated in [1]. We shows that for a function $\eta$ satisfying the Rotation number hypothesis (H), for every $s\in B$, the continuous solutions $\psi_{\eta,s}(\cdot)$ and $\psi_{\eta,1-\overline{s}}(\cdot)$ of the differential equation (2) cannot both be bounded on $1,+\infty)$.
Theorem 4.
Let $\eta\in L^{\infty}(\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). Suppose that the function $\mu_{\eta}$ satisfies:
$\forall\beta\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,\beta)\neq 0.$
Then, for every $s\in B$, we have
$\big(\mu_{\eta}(s),\,\mu_{\eta}(1-\overline{s})\big)\neq(0,0).$
2 Main proposition
Let us introduce the following notation.
Notation 5.
We denote by $I\subset\mathbb{C}$ the following set
$I:=\Big\{w\in\mathbb{C}:\Re(w)\in[0,1],\ \Re(w)\neq\tfrac{1}{2},\ \Im(w)\neq 0\Big\}.$
Consider the continuous solutions of the differential equation (2) as given in Notation 1. For every $w\in I$, we denote by $\delta_{w}$ the function
$\delta_{\eta,w}(t):=\frac{1}{2\Re(w)-1}\Big(\psi_{\eta,w}(t)-\psi_{\eta,1-\overline{w}}(t)\Big),\quad\forall t\geq 1.$
The differential equation (2) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces an order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We start to exhibit this order between solutions through the following proposition.
Proposition 6.
For every $s:=\sigma+i\tau\in I$, the function $\delta_{\eta,s}$, defined in Notation (5), satisfies the following integral equation:
$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
Proof.
Let $w\in\mathbb{C}_{+}$ and denote $w:=\alpha+i\,\beta$. For every $t\geq 1$ denote
$x_{w}(t):=t^{-\frac{1}{2}}\psi_{\eta,w}(t).$ (5)
According to Notation 1, we obtain the following differential equation
$\dot{x}_{w}(t)=\big(\alpha-\frac{1}{2}\big)\,t^{-1}x_{w}(t)+(1-w)\,t^{-\frac{3}{2}-i\,\beta}\,\eta(t),\quad x_{w}(1)=1.$ (6)
Let $s:=\sigma+i\,\tau\in I$. From the previous equation, we have
$\dot{x_{s}}(t)+\dot{x}_{1-\overline{s}}(t)=(\sigma-\frac{1}{2})t^{-1}\big({x_{s}}(t)-{x}_{1-\overline{s}}(t)\big)+(1-2i\tau)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
Use the fact that $x_{s}(1)=x_{1-\overline{s}}(1)=1$ and integrate, we obtain
${x_{s}}(t)+{x}_{1-\overline{s}}(t)$ $=(\sigma-\frac{1}{2})\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\overline{s}}(u)\big)\,du$
$+2+(1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du.$
From the differential equation (6), we have
$\dot{x}_{s}(t)-\dot{x}_{1-\overline{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big(x_{s}(t)+x_{1-\overline{s}}(t)\big)$
$-(2\sigma-1)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
The two previous differential equations, gives
$\dot{x}_{s}(t)-\dot{x}_{1-\overline{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{-1}\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\overline{s}}(u)\big)\,du$
$+(2\sigma-1)t^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du\Big]$
$+(2\sigma-1)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
Use the fact that ${x}_{s}(1)-{x}_{1-\overline{s}}(1)=0$ and integrate, we obtain
${x}_{s}(t)-{x}_{1-\overline{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\overline{s}}(v)\big)\,dv\,du$
$+(2\sigma-1)\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+(2\sigma-1)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u),\ du.$ (7)
By the notation (5), we have
${x_{s}}(t)-{x}_{1-\overline{s}}(t)=t^{-\frac{1}{2}}\big({\psi_{\eta,s}}(t)-{\psi}_{\eta,1-\overline{s}}(t)\big).$
By the Notation (5), we have
$\delta_{\eta,s}(t)=\frac{1}{2\sigma-1}\big({\psi_{\eta,s}}(t)-{\psi}_{\eta,1-\overline{s}}(t)\big).$
Then
${x_{s}}(t)-{x}_{1-\overline{s}}(t)=(2\sigma-1)t^{-\frac{1}{2}}\delta_{\eta,s}(t),$
In term of $\delta_{s}$ the integral equation (7) becomes
$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
3 Proof of the Theorem 4
Lemma 7.
For every $s\in I$ with $\Re(s)=\sigma$, the function $\delta_{\eta,s}$, defined in Notation (5), satisfies the following Volterra integral equation:
$\delta_{\eta,s}(t)=\delta_{\eta,1+i\Im(s)}(t)-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
Proof.
By the Proposition 6 every $s\in I$, we have
$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
By definition of the set II in Notation 5, we have in particular,
$t^{-\frac{1}{2}}\delta_{\eta,1+i\Im(s)}(t)=$ $\frac{1}{4}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\Im(s)}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
Subtracting the two previous equations gives:
$\delta_{\eta,s}(t)-\delta_{\eta,1+i\Im(s)}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$-\frac{1}{4}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\Im(s)}(v)\,dv\,du.$
After two differentiations, to the second‑order linear differential equation
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\Im(s)}(t)\Big)=\sigma(\sigma-1)t^{-2}\delta_{\eta,s}(t),$
with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\Im(s)}(1)$ and $\delta_{\eta,s}^{\prime}(1)=\delta_{\eta,1+i\Im(s)}^{\prime}(1)$. Solving this Cauchy problem gives the explicit representation
$\delta_{\eta,s}(t)=\delta_{\eta,1+i\Im(s)}(t)-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
Proof of the Theorem 4.
Let $s\in B$ and $\sigma=\Re(s)$. By the Lemma 7, we have
$\delta_{\eta,s}(t)$ $=\delta_{\eta,1+i\Im(s)}(t)$ (8)
$-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
By equation (3), we have
$\delta_{\eta,1+i,\Im(s)}(t)$ $=t\Big[1-i\ \Im(s)\int_{1}^{t}u^{-2-i\Im(s)}\eta(u)\,du\Big]$
$-1-(1-i\ \Im(s))\int_{1}^{t}u^{-1-i\Im(s)}\eta(u)\,du,\quad\forall t\geq 1,$
which can be written as
$\delta_{\eta,1+i\Im(s)}(t)$ $=-t\mu\big(1+i\Im(s)\big)+i\ \Im(s)t\int_{t}^{+\infty}u^{-2-i\Im(s)}\eta(u)\,du$
$-1-\rho\frac{1-i\ \Im(s)}{i\Im(s)}\big(1-t^{-i\Im(s)}\big)$
$-(1-i\ \Im(s))\int_{1}^{t}u^{-1-i\Im(s)}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
Applying the integration by parts formula to the last integral of the previous equation and using the hypothesis (H), there exists $M_{\Im(s)}>0$ such that
$\Big|\delta_{\eta,1+i\Im(s)}(t)+t\mu\big(1+i\Im(s)\big)\Big|<M_{\Im(s)},\quad\forall t>1.$ (9)
By hypothesis, $\mu_{\eta}(1+i\beta)\neq 0$. Assume $\Im(\mu_{\eta}(1+i\beta))\neq 0$ (the real case is analogous). Equation (9) implies that for sufficiently large $t$, $\Im(\delta_{\eta,1+i\Im(s)}(t))$ maintains a constant sign and grows linearly. It follows from (8) that:
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}\to 0^{+}\text{ as }t\to+\infty.$ (10)
In fact
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}=1-\frac{\sigma(1-\sigma)}{2\sigma-1}I(t),$
where
$I(t)=\int_{1}^{t}\left[\frac{\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}\right]v^{-1}\,\Im\big(\delta_{\eta,1+i\Im(s)}(v)\big)\,dv$
From estimate (9), as $t\to+\infty$ we have
$\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)=-\,t\Im\big(\mu_{\eta}(1+i\tau)\big)+O(1).$
Hence
$\frac{\Im\big(\delta_{\eta,1+i\tau}(v)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}=\frac{v}{t}+o(1)$
uniformly for $1\leq v\leq t$. Substituting this relation into the integral kernel gives
$I(t)\sim\int_{1}^{t}\left[t^{\sigma}v^{-\sigma}-t^{1-\sigma}v^{\sigma-1}\right]v^{-1}\frac{v}{t}\,dv.$
Thus
$I(t)\sim\frac{1}{t}\int_{1}^{t}\left(t^{\sigma}v^{-\sigma}-t^{1-\sigma}v^{\sigma-1}\right)dv.$
We compute
$\int_{1}^{t}t^{\sigma}v^{-\sigma}dv=t^{\sigma}\frac{t^{1-\sigma}-1}{1-\sigma}=\frac{t-t^{\sigma}}{1-\sigma},$
and
$\int_{1}^{t}t^{1-\sigma}v^{\sigma-1}dv=t^{1-\sigma}\frac{t^{\sigma}-1}{\sigma}=\frac{t-t^{1-\sigma}}{\sigma}.$
Therefore
$I(t)\sim\frac{1}{t}\left[\frac{t-t^{\sigma}}{1-\sigma}-\frac{t-t^{1-\sigma}}{\sigma}\right].$
After simplification we obtain
$I(t)\sim\frac{(2\sigma-1)t-\sigma t^{\sigma}+(1-\sigma)t^{1-\sigma}}{\sigma(1-\sigma)t}.$
Substituting into the previous expression yields
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}\sim 1-\frac{\sigma(1-\sigma)}{2\sigma-1}\frac{(2\sigma-1)t-\sigma t^{\sigma}+(1-\sigma)t^{1-\sigma}}{\sigma(1-\sigma)t}.$
After simplification we obtain
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}\sim\frac{\sigma}{(2\sigma-1)t^{1-\sigma}}-\frac{1-\sigma}{(2\sigma-1)t^{\sigma}}>0\quad\text{as}\quad t\to+\infty.$
If $\mu_{\eta}(s)=\mu_{\eta}(1-\overline{s})=0$, Lemma 3 implies that there exists $C_{s}>0$ such that
$\left|\Im\big(\delta_{\eta,s}(t)\big)-\Im\left(\rho\left(\frac{s-1}{s}+\frac{\overline{s}}{1-\overline{s}}\right)t^{-i\Im(s)}\right)\right|<C_{s}t^{-1},\quad\forall t\geq 1.$
Since $s\in B$, the $\Re(s)\neq\frac{1}{2}$ and $\Im(s)\neq 0$. Use the fact that $\rho\neq 0$, then
$\rho\big(\frac{s-1}{s}+\frac{\overline{s}}{1-\overline{s}}\big)\neq 0.$
Hence, there exists a sequence $(t_{n})_{n\in\mathbb{N}}\subset[1,+\infty)$ such that $\lim_{n\to+\infty}t_{n}=+\infty$ and such that
$\Im\big(\delta_{\eta,s}(t_{n})\big)=0,\quad\forall n\in\mathbb{N},$
From equation (9) and the assumption $\Im(\mu_{\eta}(1+i\beta))\neq 0$, there exists $n_{s}\in\mathbb{N}$ such that
$\frac{\Im\big(\delta_{\eta,s}(t_{n})\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t_{n})\big)}=0,\quad\forall n\geq n_{s},$
In other words, the function
$t\mapsto\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}$
oscillates like $t^{-1}\Im\big(\omega_{s}t^{-i\Im(s)}\big)$ with $\omega_{s}\in\mathbb{C}^{*}$, preventing the ratio in (10) from remaining strictly positive or converging as required, thus yielding a contradiction. ∎

### 15 Apr 2026 v3

@@
 We identify a Rotation number hypothesis on the non-homogeneous term, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\overline{s}$ lying in the critical strip, are both bounded on $1,+\infty)$, then $\Re(s)=\tfrac{1}{2}$.
-Keywords: Non-homogeneous complex linear differential equation, bounded solutions, Volterra integral equation.
+Keywords: Non-homogeneous complex linear differential equation, bounded solutions, Euler differential equation, Volterra integral equation.
 MSC subject classifications: 34A30, 34E05
-1 Main result
+1 Introduction
 In this manuscript, we study the following non-homogeneous linear complex differential equations:
 $\dot{\theta}=wt^{-1}\theta+t^{-1}\eta(t),\quad\theta(1)=\frac{1}{1-w},\quad\theta:[1,+\infty)\to\mathbb{C},$ (1)
-where $w\in\mathbb{C}$ such that $\Im(w)\neq 0$ is the parameter and the non-homogeneous term $\eta\in L^{\infty}(\mathbb{C})$ where $L^{\infty}(\mathbb{C})$ is the set of bounded and locally integrable functions from $1,+\infty)$ to $\mathbb{C}$.
-As formulated in the Dynamical Conjecture in [1], we assume that the function $\eta$ satisfies the following Rotation number hypothesis:
-$\exists\rho\in\mathbb{C}^{*}:\quad\sup_{t\geq 1}\Big|\int_{1}^{t}\,\big(\eta(u)-\rho\big)du\Big|<+\infty,$ (H)
-The complex number $\rho$ is called the rotation number of $\eta$.
+where $w\in\mathbb{C}$ such that $\Im(w)\neq 0$ is the parameter and the non-homogeneous term $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ where $\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ is the set of bounded and locally integrable functions from $1,+\infty)$ to $\mathbb{C}$.
 We are interested in the initial conditions of the bounded solutions of the previous differential equation.
@@
 $t\in[1,+\infty),\quad x(1)=1,\quad x:[1,+\infty)\to\mathbb{C}.$
-Since $\eta\in L^{\infty}(\mathbb{C})$, the function
+Since $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, the function
 $t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1.$
@@
 Notation 1.
-For every $\eta\in L^{\infty}(\mathbb{C})$ and $w\in\mathbb{C}$, we denote by $\psi_{\eta,w}(\cdot)$ the unique continuous solution of the differential equation (2) given by equation (3).
+For every $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and $w\in\mathbb{C}$, we denote by $\psi_{\eta,w}(\cdot)$ the unique continuous solution of the differential equation (2) given by equation (3).
 Notation 2.
@@
 $\mathbb{C}_{+}:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
-For every $\eta\in L^{\infty}(\mathbb{C})$, let $\mu_{\eta}$ denote the function $\mu_{\eta}:\mathbb{C}_{+}\to\mathbb{C}$, defined as
-$\mu_{\eta}(w)=-1-(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall w\in\mathbb{C}_{+}.$
-The function $\mu_{\eta}$ is defined for all $s\in\mathbb{C}_{+}$.
-Indeed, the integral is absolutely convergent, since $\eta\in L^{\infty}(\mathbb{C})$ and since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$.
+For every $f\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, let $\mu_{f}$ denote the function $\mu_{f}:\mathbb{C}_{+}\to\mathbb{C}$, defined as
+$\mu_{f}(w)=-1-(1-w)\int_{1}^{+\infty}u^{-1-w}f(u)\,du,\quad\forall w\in\mathbb{C}_{+}.$
+The function $\mu_{f}$ is defined for all $s\in\mathbb{C}_{+}$.
+Indeed, the integral is absolutely convergent, since $f\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$.
 In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
 Lemma 3.
-Let $\eta\in L^{\infty}(\mathbb{C})$ and $w\in\mathbb{C}_{+}$. Consider the continuous solution $\psi_{\eta,w}(\cdot)$ of the differential equation (2), as introduced in Notation 1. Then,
+Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and $w\in\mathbb{C}_{+}$.
+Consider the continuous solution $\psi_{\eta,w}(\cdot)$ of the differential equation (2), as introduced in Notation 1.
+Then,
 $\sup_{t\geq 1}\left|\psi_{\eta,w}(t)\right|<+\infty\quad\text{if and only if}\quad\mu_{\eta}(w)=0,$
-where the function $\mu_{\eta}$ is defined in Notation 2. Furthermore, if $\eta$ satisfies the rotation number hypothesis (H), then
-$\sup_{t\geq 1}\left|t\left(\psi_{\eta,w}\big(t\big)-\rho\frac{w-1}{w}t^{-i\Im(w)}\right)\right|<+\infty,$
-where $\rho$ denotes the rotation number of $\eta$.
+where the function $\mu_{\eta}$ is defined in Notation 2.
 Proof.
@@
 $\psi_{\eta,w}(t)=-(1-w)t^{\Re(w)}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (4)
-By Hypothesis $\eta\in L^{\infty}(\mathbb{C})$ then there exists a constant $c>0$ such that $|\eta(t)|<c$ almost everywhere, we obtain
+By Hypothesis $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ then there exists a constant $c>0$ such that $|\eta(t)|<c$ for all $t\geq 1$, we obtain
 $\sup_{t\geq 1}|\psi_{\eta,w}(t)|\leq c\frac{|1-w|}{\Re(w)}<+\infty.$
-Now, prove the second item of the present Lemma. Suppose that the function $\eta$ satisfies the hypothesis (H). Then there exists $\tilde{c}>0$ such that
-$\Big|\int_{1}^{t}\,\big(\eta(u)-\rho\big)du\Big|<\tilde{c},\quad\forall t\geq 1,$
-where $\rho$ is the rotation number of $\eta$. Hence, the equation (4) can be written as
-$\psi_{\eta,w}(t)=\rho\frac{w-1}{w}t^{-i\,\Im(w)}-(1-w)t^{{\Re(w)}}\int_{t}^{+\infty}u^{-1-w}(\eta(u)-\rho)du,$
-Using the integration by parts formula, we obtain
-$\psi_{\eta,w}(t)=\rho\frac{w-1}{w}t^{-i\,\Im(w)}-(1-w^{2})t^{\Re(w)}\int_{t}^{+\infty}u^{-2-w}\int_{t}^{u}(\eta(v)-\rho)dvdu.$
-For every $t\geq 1$, we get
-$\left|\psi_{\eta,w}(t)-\rho\frac{w-1}{w}t^{-i\,\Im(w)}\right|<2\tilde{c}\frac{|1-w^{2}|}{1+\Re(w)}t^{-1}.$
+2 Main result
+As formulated in the Dynamical Conjecture in [1], we assume that the function $\eta$ satisfies the following Rotation number hypothesis:
+$\exists\rho_{\eta}\in\mathbb{C}^{*}:\quad\sup_{t\geq 1}\Big|\int_{1}^{t}\,\big(\eta(u)-\rho_{\eta}\big)du\Big|<+\infty,$ (H)
+The complex number $\rho_{\eta}$ is called the rotation number of $\eta$.
+Notice that if $\rho_{\eta}$ exists, then it is unique; this can be proved by contradiction.
+Denote by $\mathcal{L}[\eta]$ the following transformation
+$\mathcal{L}[\eta](t):=\int_{1}^{t}\,\big(\eta(u)-\rho_{\eta}\big)du,\quad\forall t\geq 1.$
+The Rotation number hypothesis implies that $\mathcal{L}[\eta]\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$.
+Hence, the function $\mu_{\mathcal{L}[\eta]}$ is defined on $\mathbb{C}_{+}$ as given in the Notation 2.
 We denote by $B\subset\mathbb{C}$ the critical strip, excluding the critical line and the real axis, defined as
@@
 Theorem 4.
-Let $\eta\in L^{\infty}(\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). Suppose that the function $\mu_{\eta}$ satisfies:
-$\forall\beta\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,\beta)\neq 0.$
+Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). Suppose that the functions $\mu_{\eta}$ and $\mathcal{L}[\eta]$ satisfies:
+$\forall x\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,x)\neq 0\quad\text{and}\quad\mu_{\mathcal{L}[\eta]}(1+i\,x)\neq\frac{\rho_{\eta}-ix}{1+ix}-\frac{1}{|1+i\,x|^{2}}.$
 Then, for every $s\in B$, we have
 $\big(\mu_{\eta}(s),\,\mu_{\eta}(1-\overline{s})\big)\neq(0,0).$
+Remark 5.
+In the proof of the Lemma 9, we note that
+$\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big]=\lim_{r\to_{>}0}\mu_{\eta}(r+i\,x),$
+which is the analytic extension of the function $\mu_{\eta}$ to the line $\Re(z)=0$. Consequently, the condition of the previous theorem is equivalent to
+$\forall x\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,x)\neq 0\quad\text{and}\quad\mu_{\eta}(i\,x)\neq 0.$
 3 Main proposition
 Let us introduce the following notation.
-Notation 5.
+Notation 6.
 We denote by $I\subset\mathbb{C}$ the following set
 $I:=\Big\{w\in\mathbb{C}:\Re(w)\in[0,1],\ \Re(w)\neq\tfrac{1}{2},\ \Im(w)\neq 0\Big\}.$
-Consider the continuous solutions of the differential equation (2) as given in Notation 1. For every $w\in I$, we denote by $\delta_{w}$ the function
+Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$.
+Consider the continuous solutions of the differential equation (2) as given in Notation 1.
+For every $w\in I$, we denote by $\delta_{w}$ the function
 $\delta_{\eta,w}(t):=\frac{1}{2\Re(w)-1}\Big(\psi_{\eta,w}(t)-\psi_{\eta,1-\overline{w}}(t)\Big),\quad\forall t\geq 1.$
@@
 Proposition 7.
-For every $s:=\sigma+i\tau\in I$, the function $\delta_{\eta,s}$, defined in Notation (5), satisfies the following integral equation:
-$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
-$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
-$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
+Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$.
+For every $s:=\sigma+i\tau\in I$, the functions $\delta_{\eta,s}$ and $\delta_{\eta,1+i\tau}$, defined in Notation (6), satisfies the following Volterra integral equation:
+$\delta_{\eta,1+i\tau}(t)$ $=\sigma(1-\sigma)t^{1-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\delta_{\eta,1+i\tau}(v)\,dv\ du$
+$+\delta_{\eta,s}(t),\quad\forall t\geq 1,$
+Lemma 8.
+Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$.
+For every $s:=\sigma+i\tau\in I$, the functions $\delta_{\eta,s}$ and $\delta_{\eta,1+i\tau}$, defined in Notation (6), satisfies the following Euler differential equation:
+$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),\quad\forall t\geq 1,$
+with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$.
 Proof.
@@
 $+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
-3 Proof of the Theorem 4
-Lemma 7.
-For every $s\in I$ with $\Re(s)=\sigma$, the function $\delta_{\eta,s}$, defined in Notation (5), satisfies the following Volterra integral equation:
-$\delta_{\eta,s}(t)=\delta_{\eta,1+i\Im(s)}(t)-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
-Proof.
-By the Proposition 6 every $s\in I$, we have
-$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
-$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
-$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
-By definition of the set II in Notation 5, we have in particular,
-$t^{-\frac{1}{2}}\delta_{\eta,1+i\Im(s)}(t)=$ $\frac{1}{4}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\Im(s)}(v)\,dv\,du$
+By definition of the set II in Notation 6, we have in particular,
+$t^{-\frac{1}{2}}\delta_{\eta,1+i\tau}(t)=$ $\frac{1}{4}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\tau}(v)\,dv\,du$
 $+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
@@
 Subtracting the two previous equations gives:
-$\delta_{\eta,s}(t)-\delta_{\eta,1+i\Im(s)}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
-$-\frac{1}{4}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\Im(s)}(v)\,dv\,du.$
+$\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
+$-\frac{1}{4}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\tau}(v)\,dv\,du.$
 After two differentiations, to the second‑order linear differential equation
-$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\Im(s)}(t)\Big)=\sigma(\sigma-1)t^{-2}\delta_{\eta,s}(t),$
-with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\Im(s)}(1)$ and $\delta_{\eta,s}^{\prime}(1)=\delta_{\eta,1+i\Im(s)}^{\prime}(1)$.
-Solving this Cauchy problem gives the explicit representation
-$\delta_{\eta,s}(t)=\delta_{\eta,1+i\Im(s)}(t)-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
+$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),$
+with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$. ∎
+Proof of the Proposition 7.
+By the Lemma 8 every $s:=\sigma+i\,\tau\in I$, we have
+$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),$
+which can be written as
+$t^{1-\sigma}\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-1-\sigma}\delta_{\eta,s}(t),$
+Use the fact that $\delta_{\eta,s}(1)-\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$ and use the integration by part formula,
+$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)\int_{1}^{t}\,u^{-\sigma}\frac{d}{du}\big(\delta_{\eta,s}(u)-\delta_{\eta,1+i\tau}(u)\big)\,du$
+$=-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,s}(u)\,du,$
+$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)t^{-\sigma}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$
+$-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\big(\delta_{\eta,s}(u)-\delta_{\eta,1+i\tau}(u)\big)\,du$
+$=-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,s}(u)\,du,$
+We obtained
+$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)t^{-\sigma}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$
+$+\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,1+i\tau}(u)du$
+$=0,\quad\forall t\geq 1.$
+That can be written as
+$\frac{d}{dt}\Big(t^{-1+\sigma}\big(\delta_{\eta,1+i\tau}(t)-\delta_{\eta,s}(t)\big)\Big)$ $=\sigma(1-\sigma)t^{-2(1-\sigma)}\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,1+i\tau}(u)du$
+$+\frac{d}{dt}\big(t^{-1+\sigma}\delta_{\eta,s}(t)\big),\quad\forall t\geq 1,$
+Integrate
+$\delta_{\eta,1+i\tau}(t)$ $=\sigma(1-\sigma)t^{1-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\delta_{\eta,1+i\tau}(v)\,dv\ du$
+$+\delta_{\eta,s}(t),\quad\forall t\geq 1,$
+4 Proof of the Theorem 4
+Lemma 9.
+Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ satisfy the Rotation Number Hypothesis (H).
+For every $x\in\mathbb{R}^{*}$, the function $\delta_{\eta,1+ix}$, defined in Notation (6), satisfies the following equation
+$\sup_{t\geq 1}\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+ix}(u)\,du+\alpha_{x}t-\beta_{x}\ln(t)\Big|<+\infty,$
+where
+$\alpha_{x}:=\mu_{\eta}\big(1+ix\big)\ \text{and}\ \beta_{x}:=\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big].$
+Proof.
+Let $x\in\mathbb{R}^{*}$. By the Notation 6 and equation (3), we have
+$\delta_{\eta,1+i,x}(t)$ $=t\Big[1-i\,x\int_{1}^{t}u^{-2-ix}\eta(u)\,du\Big]$
+$-1-(1-i\,x)\int_{1}^{t}u^{-1-ix}\eta(u)\,du,\quad\forall t\geq 1,$
+which can be written as
+$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)+i\,xt\int_{t}^{+\infty}u^{-2-ix}\eta(u)\,du$
+$-1-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)$
+$-(1-i\,x)\int_{1}^{t}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
+Use the Rotation number hypothesis (H) and use the integration by formula
+$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)+\rho\frac{i\,x}{1+ix}t^{-ix}$
+$-1-|1+i\,x|^{2}\int_{1}^{+\infty}u^{-2-ix}\int_{1}^{u}\,\big(\eta(v)-\rho\big)\,dv\,du$
+$+i\,xt\int_{t}^{+\infty}u^{-2-ix}\big(\eta(u)-\rho\big)\,du$
+$+(1-i\,x)\int_{t}^{+\infty}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
+Equivalent to,
+$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)+\rho\frac{i\,x}{1+ix}t^{-ix}$
+$-1-|1+i\,x|^{2}\frac{1}{ix}\big(1+\mu_{\mathcal{L}[\eta]}(1+i\,x)\big)$
+$+i\,xt\int_{t}^{+\infty}u^{-2-ix}\big(\eta(u)-\rho\big)\,du$
+$+(1-i\,x)\int_{t}^{+\infty}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
+or,
+$\delta_{\eta,1+ix}(t)$ $=-t\alpha_{x}+\beta_{x}$
+$-\rho\frac{1}{ix(1+ix)}t^{-ix}$
+$+i\,x(2+ix)t\int_{t}^{+\infty}u^{-3-ix}\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\,\,du$
+$+|1+i\,x|^{2}\int_{t}^{+\infty}u^{-2-ix}\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\,du,\quad\forall t\geq 1.$
+where
+$\alpha_{x}:=\mu_{\eta}\big(1+ix\big)\ \text{and}\ \beta_{x}:=\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big].$
+Using the hypothesis (H), we have
+$\sup_{u\geq t\geq 1}\Big|\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\Big|<+\infty,$
+Then
+$\sup_{t\geq 1}\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+ix}(u)\,du+\alpha_{x}t-\beta_{x}\ln(t)\Big|<+\infty,$
 Proof of the Theorem 4.
-Let $s\in B$ and $\sigma=\Re(s)$. By the Lemma 7, we have
-$\delta_{\eta,s}(t)$ $=\delta_{\eta,1+i\Im(s)}(t)$ (8)
-$-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
-By equation (3), we have
-$\delta_{\eta,1+i,\Im(s)}(t)$ $=t\Big[1-i\ \Im(s)\int_{1}^{t}u^{-2-i\Im(s)}\eta(u)\,du\Big]$
-$-1-(1-i\ \Im(s))\int_{1}^{t}u^{-1-i\Im(s)}\eta(u)\,du,\quad\forall t\geq 1,$
-which can be written as
-$\delta_{\eta,1+i\Im(s)}(t)$ $=-t\mu\big(1+i\Im(s)\big)+i\ \Im(s)t\int_{t}^{+\infty}u^{-2-i\Im(s)}\eta(u)\,du$
-$-1-\rho\frac{1-i\ \Im(s)}{i\Im(s)}\big(1-t^{-i\Im(s)}\big)$
-$-(1-i\ \Im(s))\int_{1}^{t}u^{-1-i\Im(s)}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
-Applying the integration by parts formula to the last integral of the previous equation and using the hypothesis (H), there exists $M_{\Im(s)}>0$ such that
-$\Big|\delta_{\eta,1+i\Im(s)}(t)+t\mu\big(1+i\Im(s)\big)\Big|<M_{\Im(s)},\quad\forall t>1.$ (9)
-By hypothesis, $\mu_{\eta}(1+i\beta)\neq 0$.
-Assume $\Im(\mu_{\eta}(1+i\beta))\neq 0$ (the real case is analogous).
-Equation (9) implies that for sufficiently large $t$, $\Im(\delta_{\eta,1+i\Im(s)}(t))$ maintains a constant sign and grows linearly.
-It follows from (8) that:
-$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}\to 0^{+}\text{ as }t\to+\infty.$ (10)
-In fact
-$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}=1-\frac{\sigma(1-\sigma)}{2\sigma-1}I(t),$
+Let $s:=\sigma+i\,\tau\in B$. By the Lemma 9, there exists $M_{\tau}>0$ such that
+$\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+i\tau}(u)\,du+\alpha_{\tau}t-\beta_{\tau}\ln(t)\Big|<M_{\tau},\quad\forall t\geq 1,$ (8)
+where by the hypothesis of the present Theorem we have $\alpha_{\tau}\neq 0$ and $\beta_{\tau}\neq 0$. Then there exists $\omega_{\tau}\in\mathbb{C}^{*}$ such that
+$\Re\big(\omega_{\tau}\alpha_{\tau})\Re\big(\omega_{\tau}\beta_{\tau})>0\quad\text{or}\quad\Im\big(\omega_{\tau}\alpha_{\tau})\Im\big(\omega_{\tau}\beta_{\tau})>0.$
+Suppose that
+$\Re\big(\omega_{\tau}\alpha_{\tau})\Re\big(\omega_{\tau}\beta_{\tau})>0.$ (9)
+The proof is analogous for the imaginary part. Fix $\omega_{\tau}$ satisfying the previous equation and define
+$\tilde{\delta}_{1+i\tau}(t):=\Re\big(\omega_{\tau}\phi_{\eta,\tau}(t)\big)\quad\text{and}\quad\tilde{\delta}(t):=\Re\big(\omega_{\tau}\delta_{\eta,s}(t)\big).$ (10)
+Equation (8) implies
+$\Big|\int_{1}^{t}\,u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du+\Re\big(\omega_{\tau}\alpha_{\tau})t-\Re\big(\omega_{\tau}\beta_{\tau})\,\ln(t)\Big|<M_{\tau}\,|\omega_{\tau}|,\quad\forall t\geq 1,$ (11)
+For every $c>1$, denote
+$\mathcal{A}_{c}(t):=\frac{1}{2}+\frac{1}{\pi}\arctan(t-c);\quad\forall t\geq 1.$
+For every $c>1$ and $g\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, consider the following transformation
+$\Phi_{c}[g](t):=c\,t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}g(u)\,du,\quad\forall t\geq 1.$
+Since $\lim_{t\to+\infty}\mathcal{A}_{c}(t)=1$ for every $c>1$, by equation (11), we have
+$\forall c>1:\quad\lim_{t\to+\infty}\Big|t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|=\big|\Re\big(\omega_{\tau}\alpha_{\tau})\big|>0.$ (12)
+Using equation (9), there exist a sufficiently large $c_{s}>1$ and $t_{c_{s}}>1$ such that for all $c>c_{s}$ and $t>t_{c_{s}}$ we have
+$\Big|t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|=\max_{z\in[1,t]}\Big|z^{-1}\int_{1}^{z}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|,$ (13)
+Implies
+$\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|=\max_{z\in[1,t]}\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](z)\Big|,\quad\forall c>c_{s},\quad\forall t>t_{c_{s}}.$ (14)
+The purpose of introducing $\mathcal{A}_{c}$ is to obtain the previous equality.
+It ensures that the peaks of the function $t\mapsto\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|$ are bounded above by the limiting value which is equal to $c|\Re\big(\omega_{\tau}\alpha_{\tau})|$.
+By the Proposition 7, for all $t\geq 1$ we have
+$\frac{t^{-1}}{\sigma(1-\sigma)}\Big(\tilde{\delta}_{1+i\tau}(t)$ $-\tilde{\delta}_{s}(t)\Big)$
+$=t^{-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\tilde{\delta}_{1+i\tau}(v)\,dv\ du,$
+Applying the transformation $\Phi_{c}$,
+$\frac{t}{\sigma(1-\sigma)}\Big(\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)-\Phi_{c}[\tilde{\delta}_{s}](t)\Big)$
+$=c\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}\,\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\mathcal{A}_{c}(z)z^{-1}\tilde{\delta}_{1+i\tau}(z)\,dz\,dv\ du,$
+Use the integration by part formula,
+$\frac{t}{\sigma(1-\sigma)}\Big(\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)-\Phi_{c}[\tilde{\delta}_{s}](t)\Big)$
+$=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\Phi_{c}[\tilde{\delta}_{1+i\tau}](v)\,dv\,du$
+$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\Phi_{c}[\tilde{\delta}_{1+i\tau}](z)\,dz\,dv\,du,$
+We have
+$\frac{4}{\pi}\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}=-\frac{\sigma\,z^{-\sigma-1}\left(\frac{\pi}{2}+\arctan(z-c)\right)+\frac{z^{-\sigma}}{1+(z-c)^{2}}}{\left(\frac{\pi}{2}+\arctan(z-c)\right)^{2}}<0,\quad\forall z\geq 1.$
+By equation (14), for all $c>c_{s}$ and for all $t\geq t_{c_{s}}$ , we obtain
+$t\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|<t\Big|\Phi_{c}[\tilde{\delta}_{s}](t)\Big|+\sigma(1-\sigma)\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|K(\sigma,c,t),$ (15)
 where
-$I(t)=\int_{1}^{t}\left[\frac{\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}\right]v^{-1}\,\Im\big(\delta_{\eta,1+i\Im(s)}(v)\big)\,dv$
-From estimate (9), as $t\to+\infty$ we have
-$\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)=-\,t\Im\big(\mu_{\eta}(1+i\tau)\big)+O(1).$
-Hence
-$\frac{\Im\big(\delta_{\eta,1+i\tau}(v)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}=\frac{v}{t}+o(1)$
-uniformly for $1\leq v\leq t$. Substituting this relation into the integral kernel gives
-$I(t)\sim\int_{1}^{t}\left[t^{\sigma}v^{-\sigma}-t^{1-\sigma}v^{\sigma-1}\right]v^{-1}\frac{v}{t}\,dv.$
-Thus
-$I(t)\sim\frac{1}{t}\int_{1}^{t}\left(t^{\sigma}v^{-\sigma}-t^{1-\sigma}v^{\sigma-1}\right)dv.$
-We compute
-$\int_{1}^{t}t^{\sigma}v^{-\sigma}dv=t^{\sigma}\frac{t^{1-\sigma}-1}{1-\sigma}=\frac{t-t^{\sigma}}{1-\sigma},$
+$K(\sigma,c,t)$ $:=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv\,du$
+$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\,dz\,dv\,du,\quad\forall t\geq 1.$
+We prove the theorem by contradiction.
+Suppose that $\mu_{\eta}(s)=\mu_{\eta}(1-\overline{s})=0$.
+Using the notation of $\delta_{\eta,s}$ given in Notation 6,it follows from the second item of the Lemma 3 that there exists a constant $C_{s}>0$ such that such that
+$\left|\delta_{\eta,s}(t)\right|<C_{s},\quad\forall t\geq 1.$
+Using the following compact convergence
+$\lim_{c\to+\infty}c\,\mathcal{A}_{c}(u)=\frac{1}{\pi},$
+We obtain
+$\lim_{c\to+\infty}\Big|t\Phi_{c}[{\delta}_{\eta,s}](t)\Big|=\lim_{c\to+\infty}\Big|c\int_{1}^{t}\,\mathcal{A}_{c}(u)u^{-1}\delta_{\eta,s}(u)\,du\Big|$ $\leq\frac{1}{\pi}\,C_{s}\,\ln(t),\quad\forall t\geq 1.$
+By the notation of $\tilde{\delta}_{s}$ in equation (10), we get
+$\lim_{c\to+\infty}\Big|t\Phi_{c}[\tilde{\delta}_{s}](t)\Big|<\frac{1}{\pi}|\omega_{\tau}|\,C_{s}\,\ln(t),\quad\forall t\geq 1.$
+Equation (15) gives
+$\lim_{c\to+\infty}\frac{\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|}{\ln(t)}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)<\frac{1}{\pi}|\omega_{\tau}|\,C_{s},\quad\quad\forall t\geq t_{c_{s}}.$ (16)
+Now, let us examine the asymptotic behavior of $K(\sigma,c,t)$. We recall the definition
+$\mathcal{A}_{c}(t)=\frac{1}{2}+\frac{1}{\pi}\arctan(t-c),\qquad t\geq 1,$
 and
-$\int_{1}^{t}t^{1-\sigma}v^{\sigma-1}dv=t^{1-\sigma}\frac{t^{\sigma}-1}{\sigma}=\frac{t-t^{1-\sigma}}{\sigma}.$
-Therefore
-$I(t)\sim\frac{1}{t}\left[\frac{t-t^{\sigma}}{1-\sigma}-\frac{t-t^{1-\sigma}}{\sigma}\right].$
-After simplification we obtain
-$I(t)\sim\frac{(2\sigma-1)t-\sigma t^{\sigma}+(1-\sigma)t^{1-\sigma}}{\sigma(1-\sigma)t}.$
-Substituting into the previous expression yields
-$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}\sim 1-\frac{\sigma(1-\sigma)}{2\sigma-1}\frac{(2\sigma-1)t-\sigma t^{\sigma}+(1-\sigma)t^{1-\sigma}}{\sigma(1-\sigma)t}.$
-After simplification we obtain
-$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}\sim\frac{\sigma}{(2\sigma-1)t^{1-\sigma}}-\frac{1-\sigma}{(2\sigma-1)t^{\sigma}}>0\quad\text{as}\quad t\to+\infty.$
-If $\mu_{\eta}(s)=\mu_{\eta}(1-\overline{s})=0$, Lemma 3 implies that there exists $C_{s}>0$ such that
-$\left|\Im\big(\delta_{\eta,s}(t)\big)-\Im\left(\rho\left(\frac{s-1}{s}+\frac{\overline{s}}{1-\overline{s}}\right)t^{-i\Im(s)}\right)\right|<C_{s}t^{-1},\quad\forall t\geq 1.$
-Since $s\in B$, the $\Re(s)\neq\frac{1}{2}$ and $\Im(s)\neq 0$. Use the fact that $\rho\neq 0$, then
-$\rho\big(\frac{s-1}{s}+\frac{\overline{s}}{1-\overline{s}}\big)\neq 0.$
-Hence, there exists a sequence $(t_{n})_{n\in\mathbb{N}}\subset[1,+\infty)$ such that $\lim_{n\to+\infty}t_{n}=+\infty$ and such that
-$\Im\big(\delta_{\eta,s}(t_{n})\big)=0,\quad\forall n\in\mathbb{N},$
-From equation (9) and the assumption $\Im(\mu_{\eta}(1+i\beta))\neq 0$, there exists $n_{s}\in\mathbb{N}$ such that
-$\frac{\Im\big(\delta_{\eta,s}(t_{n})\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t_{n})\big)}=0,\quad\forall n\geq n_{s},$
-In other words, the function
-$t\mapsto\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}$
-oscillates like $t^{-1}\Im\big(\omega_{s}t^{-i\Im(s)}\big)$ with $\omega_{s}\in\mathbb{C}^{*}$, preventing the ratio in (10) from remaining strictly positive or converging as required, thus yielding a contradiction.
-∎
+$K(\sigma,c,t)$ $=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv\,du$
+$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\,dz\,dv\,du,\quad\forall t\geq 1.$
+Use the integration by parts formula to obtain
+$K(\sigma,c,t)=\frac{1}{2\sigma-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}W(u)\,du,$ (17)
+with
+$W(u):=\dfrac{1}{\mathcal{A}_{c}(1)}(u^{-1+2\sigma}-1)+u^{-1+2\sigma}\int_{1}^{u}\dfrac{v^{-\sigma}}{\mathcal{A}_{c}(v)}\,dv-\int_{1}^{u}\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv.$
+Use the integration by part formula, to get
+$\mathcal{A}_{c}(u)u^{-\sigma}W(u)$ $=\frac{1}{1-\sigma}-\frac{1}{\sigma}$
+$-\frac{\mathcal{A}_{c}(u)}{1-\sigma}u^{-1+\sigma}\Big[\dfrac{\sigma}{\mathcal{A}_{c}(1)}+\int_{1}^{u}v^{1-\sigma}\frac{d}{dv}\dfrac{1}{\mathcal{A}_{c}(v)}\,dv\Big]$
+$+\frac{\mathcal{A}_{c}(u)}{\sigma}u^{-\sigma}\Big[\dfrac{1-\sigma}{\mathcal{A}_{c}(1)}+\int_{1}^{u}v^{\sigma}\frac{d}{dv}\dfrac{1}{\mathcal{A}_{c}(v)}\,dv\Big]$
+Since
+$\frac{d}{du}\mathcal{A}_{c}(u)=\frac{1}{\pi}\frac{1}{1+(u-c)^{2}},\quad\forall u\geq 1,$
+It follows that
+$\mathcal{A}_{c}(u)u^{-\sigma}W(u)$ $=\frac{1}{1-\sigma}-\frac{1}{\sigma}$
+$-\frac{\mathcal{A}_{c}(u)}{1-\sigma}u^{-1+\sigma}\Big[\dfrac{\sigma}{\mathcal{A}_{c}(1)}-\frac{1}{\pi}\int_{1}^{u}\dfrac{1}{(\mathcal{A}_{c}(v))^{2}}\frac{v^{1-\sigma}}{1+(v-c)^{2}}\,dv\Big]$
+$+\frac{\mathcal{A}_{c}(u)}{\sigma}u^{-\sigma}\Big[\dfrac{1-\sigma}{\mathcal{A}_{c}(1)}-\frac{1}{\pi}\int_{1}^{u}\dfrac{1}{(\mathcal{A}_{c}(v))^{2}}\frac{v^{\sigma}}{1+(v-c)^{2}}\,dv\Big]$
+Then
+$\lim_{c\to+\infty}\mathcal{A}_{c}(u)u^{-\sigma}W(u)=\frac{2\sigma-1}{\sigma(1-\sigma)}-\frac{\sigma}{1-\sigma}u^{-1+\sigma}+\frac{1-\sigma}{\sigma}u^{-\sigma}$
+Inserting these into the integral for $K(\sigma,c,t)$ given by equation (17), we obtain the following equation
+$\lim_{c\to+\infty}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)=\frac{1}{2\sigma-1}\,\Big(\sigma t^{\sigma}-(1-\sigma)t^{1-\sigma}\Big).$
+Using the following compact convergence
+$\lim_{c\to+\infty}c\,\mathcal{A}_{c}(u)=\frac{1}{\pi},$
+Thanks to equation (11), we get
+$\lim_{t\to+\infty}\lim_{c\to+\infty}\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|=\frac{1}{\pi}\big|\Re\big(\omega_{\tau}\alpha_{\tau})\big|>0.$
+Since $\sigma\in(0,1)\setminus\frac{1}{2}$, we obtain
+$\lim_{t\to+\infty}\lim_{c\to+\infty}\frac{\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|}{\ln(t)}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)=+\infty.$
+We obtain a contradiction with equation (16). ∎